16. The Ksp value for copper (1) bromide, CuBr, is 63 x 10. What is the solubility (in water) of this compound in grams per liter? Pr= 79.90 741 -t0
Answer #1

The salt dissolves as:

CuBr <----> Cu+ + Br-

   s s

Ksp = [Cu+][Br-]

6.3*10^-9=(s)*(s)

6.3*10^-9= 1(s)^2

s = 7.937*10^-5 M

Molar mass of CuBr = 1*MM(Cu) + 1*MM(Br)

= 1*63.55 + 1*79.9

= 143.45 g/mol

Molar mass of CuBr= 143.45 g/mol

s = 7.937*10^-5 mol/L

To covert it to g/L, multiply it by molar mass

s = 7.937*10^-5 mol/L * 143.45 g/mol

s = 1.139*10^-2 g/L

Answer: 1.14*10^-2 g/L